# SparseWorld: A Flexible, Adaptive, and Efficient 4D Occupancy World Model Powered by Sparse and Dynamic Queries

### v2

@@ Abstract
Semantic occupancy has emerged as a powerful representation in world models for its ability to capture rich spatial semantics. However, most existing occupancy world models rely on static and fixed embeddings or grids, which inherently limit the flexibility of perception. Moreover, their “in-place classification” over grids exhibits a potential misalignment with the dynamic and continuous nature of real scenarios. In this paper, we propose SparseWorld, a novel 4D occupancy world model that is flexible, adaptive, and efficient, powered by sparse and dynamic queries. We propose a Range-Adaptive Perception module, in which learnable queries are modulated by the ego vehicle states and enriched with temporal-spatial associations to enable extended-range perception. To effectively capture the dynamics of the scene, we design a State-Conditioned Forecasting module, which replaces classification-based forecasting with regression-guided formulation, precisely aligning the dynamic queries with the continuity of the 4D environment. In addition, We specifically devise a Temporal-Aware Self-Scheduling training strategy to enable smooth and efficient training. Extensive experiments demonstrate that SparseWorld achieves state-of-the-art performance across perception, forecasting, and planning tasks. Comprehensive visualizations and ablation studies further validate the advantages of SparseWorld in terms of flexibility, adaptability, and efficiency. The code is available at https://github.com/MSunDYY/SparseWorld.
@@ Introduction
In recent years, vision-centric end-to-end autonomous driving, which relies on monocular or multi-view images to plan ego vehicle’s future trajectory, has attracted increasing attention in the academic and industrial community. One branch of them, occupancy-based world models (Zheng et al. 2024; Wei et al. 2024; Li et al. 2025) effectively exploit the comprehensive and accurate spatial understanding provided by semantic occupancy representations, and have demonstrated outstanding performance in planning tasks.
@@
Early occupancy world models(Zheng et al. 2024; Gu et al. 2024), as shown in Figure 1(a), individually re-encode each frame’s occupancy observations individually into embeddings, which are then fused by a future-oriented world model and subsequently decoded. These approaches, referred to as decoupled methods, separate forecasting from perception, which hinder gradient propagation and end-to-end optimization. Moreover, the encode-decode process inevitably incurs the loss of fine-grained information, no matter how carefully it is designed.
@@
Figure 1: (a) Perception-forecasting-decoupled methods suffer from gradient detachment. (b) Grid feature-based methods relying on dense and static volumes, exhibits limited flexibility and rigid “in-place classification”. (c) We adopt dynamic query representations, which facilitate continuous and coherent 4D scene forecasting and planning.
@@
Recently, a series of grid feature-based works (Li et al. 2025; Yang et al. 2025) have utilized perception-level features as intermediate representations within the world model for per-grid forecasting, as show in Figure 1(b), thus enabling end-to-end optimization. However, such static and “in-place classification” operation exhibits a potential misalignment with the continuity of ego-motion and scene dynamics, leading to temporal inconsistency, feature drift, and cumulative errors.
@@
Both of the aforementioned world models are limited by manually predefined spatial ranges, which restricts the flexibility and adaptivity of perception. In real-world driving scenarios, where vehicle speed varies drastically, dynamically adjusting the perception range is critical for accurate forecasting and planning. Moreover, dense grids incur substantial computational and memory costs, while ignoring the inherently sparse nature of the physical world.
@@
To overcome the limitations of existing world models, and inspired by recent sparse occupancy perception models (Wang et al. 2024a; Liu et al. 2024), we propose SparseWorld, a fully sparse 4D occupancy world model built upon sparse and dynamic queries, as show in Figure 1(c). SparseWorld follows a “perceive-then-forecast” paradigm: it first adaptively constructs the extended-range occupancy queries at the current timestamp, and then regresses the future motion of scene elements relative to the ego vehicle, moving beyond traditional classification on static grids.
@@
Taking learnable queries as input, we propose a Range-Adaptive Perception module built with stacked decoders featuring temporal-spatial multi-head self-attention. To fully exploit the flexibility of queries and enhance the perception adaptability, we incorporate an Adaptive Scaling sub-module that encodes ego vehicle’s historical trajectory to modulate the initial distribution of queries. Building on the continuity and dynamics of queries, we further propose a State-Conditioned Forecasting module, where the ego query interacts with concurrent scene queries through spatial modulation. In this module, regression-guided migration replaces traditional “in-place classification” to capture the element motions for continuous and plausible forecasting.
@@
Additionally, we propose a targeted Temporal-Aware Self-Scheduling training strategy that implicitly partitions the timestamp distribution of queries and enables the model to autonomously learn timestamp assignments during training, which significantly improves both training efficiency and autonomy.
@@
To validate the effectiveness of SparseWorld, we conducted extensive experiments on the Occ3d-nuScenes (Tian et al. 2023) benchmark, comparing it with other state-of-the-art methods. The experimental results demonstrat that SparseWorld significantly outperforms dense models in both forecasting and planning tasks. Specifically, SparseWorld surpasses PreWorld (Li et al. 2025) by 20%–40% mIoU in future occupancy forecasting, and reduces collision rate in trajectory planning by half. Moreover, our SparseWorld achieves an approximate 7x speedup in inference compared to dense methods, greatly enhancing its practicality for real-world deployment.
@@
  • We develop a novel Temporal-Aware Self-Scheduling training strategy to promote smoother and more efficient training.
@@
Occupancy perception methods can be broadly categorized into dense and sparse paradigms according to their modeling strategies. Dense ones (Ma et al. 2024; Yu et al. 2023; Huang and Huang 2022; Wu et al. 2024; Kim et al. 2025; Wang et al. 2024b; Zhang et al. 2024b; Kim et al. 2025; Ye et al. 2024; Li et al. 2023b) typically construct BEV or volume features that are conceptually straightforward but incur significant computations and limited flexibility. In contrast, sparse approaches (Liu et al. 2024; Wang et al. 2024a; Shi et al. 2024; Li et al. 2023a) eliminate the reliance on dense representations. Recently, some weakly- and self-supervised approaches (Pan et al. 2024; Huang et al. 2024; Boeder, Gigengack, and Risse 2025; Sun et al. 2024a) has emerged to alleviate the reliance on expensive 3D annotations.
@@
World models aim to predict future scenes and plan ego-agent trajectories based on historical observations and actions (Ha and Schmidhuber 2018; Gao et al. 2023; Yang et al. 2024; Gao et al. 2024). Occupancy world models are required to simultaneously forecast future occupancy scenes and plan trajectories. Early works such as OccWorld (Zheng et al. 2024) and its variants (Wei et al. 2024; Jin et al. ; Xu et al. 2025) decouple occupancy perception and prediction by first encoding the observed scene and then forecasting autoregressively followed by re-encoding and decoding. Recent grid-based methods attempt to unify perception and prediction by constructing volumetric (Li et al. 2025) or BEV (Yang et al. 2025) features to represent the spatiotemporal world consistently.
@@
Planning-oriented end-to-end autonomous driving typically requires precise environmental perception. SP-T3 (Hu et al. 2022) and UniAD (Hu et al. 2023) represent the scene with a unified BEV representation, while VAD (Jiang et al. 2023) introduces a vectorized design. Zhang et al. (2024a) and Sun et al. (2024b) perform perception and planning sequentially based on sparse perception signals, whereas the more recent DriveTransformer (Jia et al. 2025) decouple and execute perception and planning in parallel, demonstrating superior performance in closed-loop inference. Some recent works have explored the use of anchor-based(Chen et al. 2024) and diffusion(Liao et al. 2025; Zheng et al. 2025) to encourage diverse trajectory outputs. In our work, we adopt autoregressive generation with L1 supervision for trajectory prediction for fair comparison.
@@
Here, NN denotes the number of queries and $P_{0}$ is learnable, which will be explained later. The initial queries $Q_{0}$ combined with modulated positions $P_{0}^{\prime}={\{\mathbf{p}_{i}^{\prime}\}}_{i=1}^{N}$, regardless of timestamps, are fed info $L$ stacked decoders for extended-range occupancy perception for current moment.
@@
Our stacked decoders follow a coarse-to-fine paradigm. Each query samples semantic information from multi-view, multi-scale feature maps extracted by the image backbone, followed by adaptive mixing introduced in (Liu et al. 2023, 2024). The queries are then fed into Temporal-Spatial Multi-Head Self-Attention (TS-MHSA) for spatiotemporal interaction.
@@
where $\bm{p_{i}}=(x_{i},y_{i},z_{i})$, $\tau_{i}$ is a learnable scaling factor controlling spatial bias inspired by SparseBEV (Liu et al. 2023), and $M_{ij}$ is a temporal masking term defined as:
@@
We construct an statistical matrix $M\in\mathbb{R}^{N\times(f+1)}$ that records the count of output points from each of NN queries corresponding to each timestamp over the entire dataset.
@@
Figure 3: Visualization of the current and future 3-second ground truth and forecasts. As highlighted in red boxes, PreWorld (Li et al. 2025), which relies on grid-based modeling and voxel-level classification, exhibits severe distortions in long horizon, whereas SparseWorld effectively mitigates such issues.
@@
We employ the widely adopted Occ3d-nuScenes (Tian et al. 2023) benchmark, which is built upon the nuScenes dataset (Caesar et al. 2020). It provides 700 training scenes and 150 validation scenes, each lasting 20 seconds, with labels provided every 0.5 seconds. The Occ3d-nuScenes dataset offers dense labels with a resolution of 200×200×16, comprising 17 semantic categories and one free category. Each occupancy grid cell has a size of 0.4m×0.4m×0.4m. Following established methodologies, we assess occupancy perception and forecasting using Intersection over Union (IoU) and mean IoU (mIoU) as metrics. IoU evaluates overleap considering only foreground and background, whereas mIoU computes the mean IoU across all 17 classes. L2 error and collision rate are applied as indicators for ego-vehicle trajectory planning.
@@
As shown in row 6 of Table 3, removing the Ego State Condition leads to a substantial drop of 0.51 mIoU, highlighting the critical role of ego motion in dynamic scene forecasting.
@@ Conclusion
In this paper, we presented SparseWorld, a fully sparse, flexible, adaptive, and efficient 4D occupancy world model. We introduced historical trajectory–guided adaptive-range perception and proposed an ego-conditioned forecasting module, where ego and scene queries interact in a bidirectional autoregressive manner, naturally aligning with the continuous dynamics of 4D scenes. To ensure smooth convergence, we designed a dedicated Temporal-Aware Self-Scheduling training strategy.
Extensive experiments demonstrate the superior forecasting and planning capabilities of SparseWorld. Detailed ablation studies and visualizations further reveal the underlying interpretability of SparseWorld and support our claims.
@@
The geometric reasoning capability (IoU) of SparseWorld still has substantial room for improvement. Moreover, its generalization ability to unseen scenarios remains to be further validated.
Notably, SparseWorld does not require dense occupancy annotations, and can be trained with only lidar point clouds and 2D labels. In future work, we plan to explore weakly supervised training schemes. Additionally, we aim to incorporate large language models (LLMs) to endow SparseWorld with more advanced reasoning abilities in complex scenes. We demonstrate that a limited number of sparse queries can effectively represent 4D scenarios, offering a new perspective for the autonomous driving community.

### v3

@@ Abstract
Semantic occupancy has emerged as a powerful representation in world models for its ability to capture rich spatial semantics. However, most existing occupancy world models rely on static and fixed embeddings or grids, which inherently limit the flexibility of perception. Moreover, their “in-place classification” over grids exhibits a potential misalignment with the dynamic and continuous nature of real scenarios. In this paper, we propose SparseWorld, a novel 4D occupancy world model that is flexible, adaptive, and efficient, powered by sparse and dynamic queries. We propose a Range-Adaptive Perception module, in which learnable queries are modulated by the ego vehicle states and enriched with temporal-spatial associations to enable extended-range perception. To effectively capture the dynamics of the scene, we design a State-Conditioned Forecasting module, which replaces classification-based forecasting with regression-guided formulation, precisely aligning the dynamic queries with the continuity of the 4D environment. In addition, We specifically devise a Temporal-Aware Self-Scheduling training strategy to enable smooth and efficient training. Extensive experiments demonstrate that SparseWorld achieves state-of-the-art performance across perception, forecasting, and planning tasks. Comprehensive visualizations and ablation studies further validate the advantages of SparseWorld in terms of flexibility, adaptability, and efficiency.
Codehttps://github.com/MSunDYY/SparseWorld
@@ Introduction
In recent years, vision-centric end-to-end autonomous driving, which predicts the ego vehicle’s future trajectory from monocular or multi-view images, has gained significant attention. Among them, occupancy-based world models (occworld; occllama; preworld) leverage semantic occupancy representations for rich spatial understanding and have shown superior planning performance.
@@
Early occupancy world models (occworld; dome), as shown in Fig. 1(a), independently encode each frame’s occupancy into embeddings, which are later fused and decoded by a future-oriented world model. Such decoupled designs separate forecasting from perception, hindering gradient flow and end-to-end optimization, while the repeated encode–decode process inevitably loses fine-grained information, no matter how carefully it is designed.
@@
Figure 1: (a) Perception-forecasting-decoupled methods. (b) Grid feature-based methods. (c) We adopt dynamic query representations, which facilitate continuous and coherent 4D scene forecasting and planning.
@@
Recently, a series of grid feature-based works (preworld; driveoccworld) have adopted perception features as intermediate representations for per-grid forecasting, as show in Figure 1(b), thus enabling end-to-end optimization. However, their static “in-place classification” operation misaligns with the continuity of ego-motion and scene dynamics, causing temporal inconsistency, feature drift, and cumulative errors.
@@
Both aforementioned world models rely on manually predefined spatial ranges, limiting perceptional flexibility and adaptivity. In real-world driving scenarios, where vehicle speed varies drastically, adaptive perception ranges are critical for accurate forecasting and planning. Moreover, dense grids incur high computational and memory costs while ignoring the inherently sparse nature of the physical world.
@@
To overcome these limitations, inspired by recent sparse occupancy perception models (opus; sparseocc), we propose SparseWorld, a fully sparse 4D occupancy world model built on sparse and dynamic queries, as show in Figure 1(c). Following a “perceive-then-forecast” paradigm, SparseWorld adaptively constructs extended-range occupancy queries, and regresses the future motion of scene elements relative to the ego vehicle, moving beyond static grid classification.
@@
We introduce a Range-Adaptive Perception module that takes learnable queries as input and employs stacked decoders featuring temporal-spatial fusion. To fully enhance the perception adaptability, an Adaptive Scaling sub-module encodes ego vehicle’s historical trajectory to modulate the initial distribution of queries. Leveraging the continuity and dynamics of queries, we further design a State-Conditioned Forecasting module, where the ego query interacts with scene queries via spatial modulation,and regression-guided migration replaces traditional “in-place classification” to for continuous and plausible motion forecasting.
@@
We further introduce a targeted Temporal-Aware Self-Scheduling training strategy that implicitly partitions query timestamps, allowing the model to autonomously learn timestamp assignments during training, which significantly improves training efficiency and autonomy.
@@
To validate the effectiveness of SparseWorld, we conducted extensive experiments on the Occ3d-nuScenes (occ3d) benchmark, comparing it with other state-of-the-art methods. The experimental results demonstrat that SparseWorld significantly outperforms dense models in both forecasting and planning tasks. Specifically, SparseWorld surpasses PreWorld (preworld) by 20%–40% mIoU in future occupancy forecasting, and reduces collision rate in trajectory planning by half. Moreover, our SparseWorld achieves an approximate 7x speedup in inference compared to dense methods, greatly enhancing its practicality for real-world deployment.
@@
  • We develop a novel Temporal-Aware Self-Scheduling training strategy to enable smooth and efficient training.
@@
Occupancy perception methods can be broadly categorized into dense and sparse paradigms according to their modeling strategies. Dense ones (cotr; flashocc; bevdet4d; heightocc; protoocc; panoocc; lightocc; protoocc; cvtocc; fbocc) typically construct BEV or volume features that are conceptually straightforward but incur significant computations and limited flexibility. In contrast, sparse approaches (sparseocc; opus; osp; voxformer) eliminate the reliance on dense representations. Recently, some weakly- and self-supervised approaches (renderocc; selfocc; gaussianflowocc; sun2024gsrender) has emerged to alleviate the reliance on expensive 3D annotations.
@@
World models aim to predict future scenes and plan ego-agent trajectories based on historical observations and actions (worldmodels; magicdrive; vidar; vista). Occupancy world models are required to simultaneously forecast future occupancy scenes and plan trajectories. Early works such as OccWorld (occworld) and its variants (occllama; occllm) decouple occupancy perception and prediction by first encoding the observed scene and then forecasting autoregressively followed by re-encoding and decoding. Recent grid-based methods attempt to unify perception and prediction by constructing volumetric (preworld) or BEV (driveoccworld) features to represent the spatiotemporal world consistently.
@@
Planning-oriented end-to-end autonomous driving typically requires precise environmental perception. SP-T3 (stp3) and UniAD (uniad) represent the scene with a unified BEV representation, while VAD (vad) introduces a vectorized design. sparsead and sparsedrive perform perception and planning sequentially based on sparse perception signals, whereas the more recent DriveTransformer (drivetransformer) decouple and execute perception and planning in parallel, demonstrating superior performance in closed-loop inference. Some recent works have explored the use of anchor-based(vadv2) and diffusion(diffusiondrive; diffusion-planner) to encourage diverse trajectory outputs. In our work, we adopt autoregressive generation with L1 supervision for trajectory prediction for fair comparison.
@@
Here, $N$ denotes the number of queries and $P_{0}$ is learnable, which will be explained later. The initial queries $Q_{0}$ combined with modulated positions $P_{0}^{\prime}={\{\mathbf{p}_{i}^{\prime}\}}_{i=1}^{N}$, regardless of timestamps, are fed info $L$ stacked decoders for extended-range occupancy perception for current moment.
@@
Our stacked decoders follow a coarse-to-fine paradigm. Each query samples semantic information from multi-view, multi-scale feature maps extracted by the image backbone, followed by adaptive mixing introduced in (sparsebev; sparseocc). The queries are then fed into Temporal-Spatial Multi-Head Self-Attention (TS-MHSA) for spatiotemporal interaction.
@@
where $\bm{p_{i}}=(x_{i},y_{i},z_{i})$, $\tau_{i}$ is a learnable scaling factor controlling spatial bias inspired by SparseBEV (sparsebev), and $M_{ij}$ is a temporal masking term defined as:
@@
We construct an statistical matrix $M\in\mathbb{R}^{N\times(f+1)}$ that records the count of output points from each of $N$ queries corresponding to each timestamp over the entire dataset.
@@
Figure 3: Visualization of the current and future 3-second ground truth and forecasts. As highlighted in red boxes, PreWorld (preworld), which relies on grid-based modeling and voxel-level classification, exhibits severe distortions in long horizon, whereas SparseWorld effectively mitigates such issues.
@@
We employ the widely adopted Occ3d-nuScenes (occ3d) benchmark, which is built upon the nuScenes dataset (nuscenes). It provides 700 training scenes and 150 validation scenes, each lasting 20 seconds, with labels provided every 0.5 seconds. The Occ3d-nuScenes dataset offers dense labels with a resolution of 200×200×16, comprising 17 semantic categories and one free category. Each occupancy grid cell has a size of 0.4m×0.4m×0.4m. Following established methodologies, we assess occupancy perception and forecasting using Intersection over Union (IoU) and mean IoU (mIoU) as metrics. IoU evaluates overleap considering only foreground and background, whereas mIoU computes the mean IoU across all 17 classes. L2 error and collision rate are applied as indicators for ego-vehicle trajectory planning.
@@
As shown in row 6 of Table 3, removing the Ego State Condition leads to a substantial drop of 0.51 mIoU, highlighting the critical role of ego motion in forecasting.
@@ Conclusion
In this paper, we proposed SparseWorld, a fully sparse, flexible, adaptive, and efficient 4D occupancy world model. Guided by ego state, SparseWorld achieves adaptive-range perception and ego-conditioned forecasting through bidirectional query interaction, aligning with the continuous dynamics of 4D scenes. A dedicated Temporal-Aware Self-Scheduling strategy further ensures stable training.
Extensive experiments demonstrate SparseWorld’s superior forecasting and planning capabilities, while ablation studies and visualizations confirm its interpretability.
@@
Despite its strong performance, SparseWorld’s geometric reasoning (IoU) and generalization to unseen scenarios remain areas for improvement. Notably, it requires only LiDAR and 2D labels without dense occupancy annotations. Future work will explore weakly supervised training and integrate large language models to enhance reasoning in complex scenes. Overall, SparseWorld demonstrates that a small number of sparse queries can effectively represent 4D scenes, offering a new perspective for autonomous driving research.
Acknowledgements
This work is funded by Xiongan AI Institute, Lenovo Research and Wuxi Research Institute of Applied Technologies, Tsinghua University under Grant 20242001120. We sincerely appreciate their supports and contributions.